# PhySense: Sensor Placement Optimization for Accurate Physics Sensing

### v2

@@ 4.2 Sea Temperature
This benchmark is derived from the GLORYS12 reanalysis dataset [20], focusing on the sea surface temperature variable. The data are downsampled to a $1^{\circ}\times 1^{\circ}°$ resolution, resulting in global ocean fields of size $360\times 180$. During training, between 10 and 100 sensors are randomly selected, with all sensors uniformly sampled from valid ocean regions, excluding land areas. The reconstruction task aims to recover dense temperature fields from these sparse observations. Our training set comprises 9,843 daily samples from 1993 to 2019, with data from 2020–2021 reserved for testing. This setup allows us to evaluate two critical aspects of model capability: (1) adaptability to varying sensor placements, and (2) generalization to temporal distribution shifts. Moreover, the presence of complex land–sea boundaries introduces significant challenges for both reconstruction and sensor placement optimization. We address this by providing a land mask as an additional input.
@@ Results
PhySense-opt consistently achieves state-of-the-art performance across all sensor numbers (average improvement 40.0%), as shown in Table 2. Besides, SSPOR [31] performs reasonably well on this benchmark, due to the relatively stable temporal variations of sea temperature over the 30-year period. This also explains its poor performance on the turbulent flow, highlighting the limitations of traditional methods in capturing complex, non-stationary physical fields.
@@
Table 3: Main results on the car aerodynamics benchmark. Due to resource constraints, only the best deterministic baseline and generative baseline are selected. To support irregular meshes, the backbone of DiffusionPDE [17] is replaced with Transolver [46].
@@
Senseiver [40] 0.1009 0.1012 0.1018 0.1022 0.1053
DiffusionPDE [17] 0.0967 0.0966 0.0987 0.1055 0.2095
@@ 4.4 Model Analysis
Beyond the main results, we explore how to best incorporate sparse sensor information. We compare our incorporation strategy with an alternative that concatenates an interpolated field—similar to the approaches used in VoroniCNN [13]. As shown in Table 4, this replacement leads to a significant performance drop, especially under low spatial coverage (e.g., 30 sensors), highlighting the advantage of our method in capturing critical information from scattered sensor observations.
Figure 6: (a) The cross-attention map visualizations from the last layer of PhySense, where sensor positions are pointed by red arrows. (b) Initial placement sensitivity analysis shows that varying initial sensor placements converge to different but equally effective placements, achieving similar reconstruction accuracy. (c) Efficiency comparison on turbulent flow: these diffusion-based methods require 1,000 generation steps (100$\times$ slower than PhySense), but achieve even worse accuracy.
@@ Appendix A Proofs of Theorems in the Method Section
where $U$ is an orthogonal matrix, i.e. $UU^{\mathsf{T}}=I$. where $U$ is an orthogonal matrix, i.e. $UU^{\mathsf{T}}=I$. All eigenvalues $\lambda_{i}(p)$ are non-negative because the covariance matrix $\Sigma_{p}$ is positive semidefinite. Since the trace is invariant under orthogonal similarity transformations, we obtain
@@
=$\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}\operatorname{tr}\!\Bigl{(}\operatorname{Var}\!\bigl{(}\mathbf{X}_{1}-\mathbf{X}_{0}\,\bigm{|}\,\mathbf{X}_{t},t,\mathbf{p},\mathbf{X}_{1}(\mathbf{p})\bigr{)}\Bigr{)}\,$
@@
[…]
@@ Appendix C Implementation Details
In this section, we provide the implementation details of our experiments, including benchmarks, metrics, and implementations. All the experiments are conducted based on PyTorch 2.1.0 [33] and on a A100 GPU server with 144 CPU cores.
@@ C.1 Benchmarks
We use a watertight 3D car mesh taken from the ShapeNet V1 [5] dataset, which contains around 50k triangular facets. The aerodynamic pressure of the car surface is simulated using OpenFOAM [19]. As shown in Fig. 7(a), we create a virtual wind tunnel to simulate the airflow around a car. The width, length and height of the wind tunnel are set to 16 meters, 30 meters and 8 meters respectively. The computational mesh around the car surface is refined to a smaller mesh size for more accurate aerodynamics modeling, which is illustrated in Fig. 7(b).
@@ C.2 Metrics
where $\mathbf{u}$ denotes the reference (ground-truth) field and $\hat{\mathbf{u}}$ represents the model reconstruction.
@@ C.3 Implementations
SSPOR
We utilize the official implementation of SSPOR, the PySensor package [7] to evaluate the reconstruction and sensor selection performance of SSPOR. SSPOR employs POD [2] to express high-dimensional data as a linear combination of several orthonormal eigenmodes obtained from the singular value decomposition (SVD). Specifically, we employ the pysensors.SSPOR class with an SVD basis to obtain the targeted number of sensors from the training data. The selected sensors are then used to reconstruct dense fields from sparse observations either with SSPOR or our method.
VoronoiCNN
VoronoiCNN integrates Voronoi tessellation with convolutional neural networks (CNNs). However, the official implementation uses a simplistic CNN that struggles with complex and dynamic fields. We replace it with a more powerful U-Net [37] architecture, using 64 base feature channels, with 4 downsampling and upsampling layers and skip connections to enhance information flow. Due to its inability to handle irregular car aerodynamic meshes, experiments are conducted only on turbulent flow and sea temperature datasets under a unified setup.
Senseiver
@@
DiffusionPDE
We borrow the model architectures of the original implementation of DiffusionPDE for the turbulent flow and sea temperature benchmarks, and employ a Transolver [46] as the backbone for the unstructured car aerodynamics benchmark. For inference, we also follow the original implementation configurations, using a total of 2,000 steps for better performance. The guidance weights we search in each benchmark are shown in Table 7.
@@ D.1 Comparsion with AdaLN-Zero
Here, we compare our sparse sensor incorporation strategy on turbulent flow with a widely used conditioning approach in diffusion models, namely the AdaLN-Zero block in DiT [34]. As shown in Table 10, replacing our design with AdaLN-Zero leads to a substantial degradation in performance, with reconstruction errors approaching 1. This is because sensor placement inherently captures localized information, whereas AdaLN-Zero aggregates all sensor data into a single global variable, which is insufficient for preserving the spatial specificity required in reconstruction tasks.
@@ D.3 Statistical Analysis
Furthermore, since our task is a reconstruction problem, and our reconstruction model essentially learns a conditional distribution, we report in the table above the mean and standard deviation of reconstruction loss under multiple samplings on the car aerodynamics benchmark, across varying numbers of sensors. As shown in Table 12 and 13, when the number of sensors exceeds 15, the standard deviation becomes relatively small, indicating stable performance. Moreover, regardless of the sensor count, once the placement is optimized, the reconstruction results exhibit consistent stability across multiple runs.

### v3

@@ 4.2 Sea Temperature
This benchmark is derived from the GLORYS12 reanalysis dataset jean2021copernicus , focusing on the sea surface temperature variable. The data are downsampled to a $1^{\circ}\times 1^{\circ}$° resolution, resulting in global ocean fields of size $360\times 180$. During training, between 10 and 100 sensors are randomly selected, with all sensors uniformly sampled from valid ocean regions, excluding land areas. The reconstruction task aims to recover dense temperature fields from these sparse observations. Our training set comprises 9,843 daily samples from 1993 to 2019, with data from 2020–2021 reserved for testing. This setup allows us to evaluate two critical aspects of model capability: (1) adaptability to varying sensor placements, and (2) generalization to temporal distribution shifts. Moreover, the presence of complex land–sea boundaries introduces significant challenges for both reconstruction and sensor placement optimization. We address this by providing a land mask as an additional input.
@@ Results
PhySense-opt consistently achieves state-of-the-art performance across all sensor numbers (average improvement 40.0%), as shown in Table 2. Besides, SSPOR manohar2018data performs reasonably well on this benchmark, due to the relatively stable temporal variations of sea temperature over the 30-year period. This also explains its poor performance on the turbulent flow, highlighting the limitations of traditional methods in capturing complex, non-stationary physical fields.
@@
Table 3: Main results on the car aerodynamics benchmark. Due to resource constraints, only the best deterministic baseline and generative baseline are selected. To support irregular meshes, the backbone of DiffusionPDE huang2024diffusionpde is replaced with Transolver wu2024transolver .
@@
Senseiver santos2023development 0.1009 0.1012 0.1018 0.1022 0.1053
DiffusionPDE huang2024diffusionpde 0.0967 0.0966 0.0987 0.1055 0.2095
@@ 4.4 Model Analysis
Beyond the main results, we explore how to best incorporate sparse sensor information. We compare our incorporation strategy with an alternative that concatenates an interpolated field—similar to the approaches used in VoroniCNN fukami2021global . As shown in Table 4, this replacement leads to a significant performance drop, especially under low spatial coverage (e.g., 30 sensors), highlighting the advantage of our method in capturing critical information from scattered sensor observations.
[…]
@@
Table 5: Ablation on progressive sensor placement optimization via a pure noise input.
Ablations Turbulent Flow
#200 #100 #50 #30
noise-inut 0.1118 0.1293 0.1630 0.2286
mix-input (origin) 0.1106 0.1257 0.1558 0.2157
Promotion 1.07% 2.78% 4.42% 5.65%
Progressive sensor placement optimization
A key advantage of our flow-based reconstructor is its inherent support for progressive sensor placement optimization. Unlike non-generative models, which operate on static inputs (e.g., pure noise or zeros), our model processes a noise-data mixture $\mathbf{X}_{t}$ throughout the flow. This design inherently embeds a spectrum of reconstruction difficulties, corresponding to different noise levels, thereby amortizing the optimization process across the entire flow. Consequently, it provides more informative, tractable feedback for sensor placement. An ablation study shown in Table 5 confirms that replacing the noise-data mix input with pure noise input in the feedback loop leads to notable performance degradation, which is more pronounced with fewer sensors, underscoring the unique benefit of this flow-process feedback.
Placement generalization
We compare our placement strategy with four more placement baselines across Senseiver and PhySense under different sensor numbers on turbulent flow dataset.
(i) Grid: We first place sensors on a uniform grid using the largest square number within the budget, and then randomly allocate the remainder. This strategy guarantees spatial coverage but remains agnostic to the physical dynamics. Practically, its performance is similar to random placement, aligning with classical sampling theory candes2006near that questions the efficacy of grid sampling for complex systems.
(ii) Min-Max: A heuristic method that promotes spread-out sensor placements by iteratively placing the next sensor at the point that maximizes the minimum distance to existing sensors. However, this approach does not yield significant improvements over random placement in our experiments.
(iii) Our Projected Gradient Descent: An upper-bound method using any reconstruction model (Senseiver or PhySense) and performs our projected gradient descent to optimize sensor placements.
(iv) Optimized Sensor from PhySense: To evaluate the generality of the learned placements by PhySense, we feed the placement into the Senseiver. The performance closely matches that of Our PGD and significantly outperforms all other baselines, particularly when the number of sensors is limited (e.g., 30), a challenging scenario to place sensors. This suggests that our learned placements effectively capture physically informative structures that generalize beyond specific reconstructors.
Table 6: Performance comparison with four additional placement baselines on the turbulent flow benchmark, measured by relative L2 error. Results obtained with limited sensors (e.g., 30) are highlighted in blue, and the top-performing methods with comparable results are shown in bold.
Placement Strategies Senseiver PhySense
#200 #100 #50 #30 #200 #100 #50 #30
Random Sample 0.1842 0.2316 0.3740 0.5746 0.1233 0.1527 0.2586 0.5176
Grid Sample 0.1615 0.2273 0.3845 0.5869 0.1235 0.1521 0.2436 0.4930
Min-Max Sample 0.1549 0.2091 0.3795 0.5747 0.1196 0.1533 0.3157 0.5022
Our PGD on reconstruction model 0.1388 0.1750 0.3025 0.4570 0.1106 0.1257 0.1558 0.2157
Optimized Sensors from PhySense 0.1533 0.1833 0.3035 0.4601 0.1106 0.1257 0.1558 0.2157
Figure 7: Visualization of information and variance distribution on turbulent flow benchmark.
Optimized sensor distribution visualization
To explore the distribution of optimized sensors, we visualize their relationship with information distribution (measured by the model’s spatial gradient magnitude) and variance distribution (point-wise variance of the physical field) on turbulent flow benchmark. Interestingly, most of the optimized sensors lie within high-variance regions. More importantly, each sensor is located near a high-information point, and their spatial distributions exhibit similar patterns. This suggests that the optimizer is not merely clustering in high-variance areas but is strategically identifying spatially separated, high-sensitivity locations which capture the field’s global dynamics. This aligns with the sensing principle that maximal information gain requires sampling diverse, non-redundant points rather than densely covering localized variance hotspots.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by the National Natural Science Foundation of China (62021002 and U2342217), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
@@ Appendix A Proofs of Theorems in the Method Section
where $U$ is an orthogonal matrix, i.e. $UU^{\mathsf{T}}=I$. where $U$ is an orthogonal matrix, i.e. $UU^{\mathsf{T}}=I$. All eigenvalues $\lambda_{i}(p)$ are non‑negative because the covariance matrix $\Sigma_{p}$ is positive semidefinite. Since the trace is invariant under orthogonal similarity transformations, we obtain
@@
$\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}\operatorname{tr}\!\Bigl(\operatorname{Var}\!\bigl(\mathbf{X}_{1}-\mathbf{X}_{0}\,\bigm|\,\mathbf{X}_{t},t,\mathbf{p},\mathbf{X}_{1}(\mathbf{p})\bigr)\Bigr)\,$
@@
$\sum_{i=1}^{d}\lambda_{i}\;\leq\;\Bigl(\sum_{i=1}^{d}\sqrt{\lambda_{i}(\mathbf{p})}\Bigr)^{2}\;\leq\;2\sum_{i=1}^{d}\lambda_{i}\quad\Longrightarrow\quad\frac{1}{2}\Bigl(\sum_{i=1}^{d}\sqrt{\lambda_{i}(\mathbf{p})}\Bigr)^{2}\leq\mathcal{L}_{\mathrm{A}}\leq\Bigl(\sum_{i=1}^{d}\sqrt{\lambda_{i}(\mathbf{p})}\Bigr)^{2}.$
@@
Remark A.1.
Furthermore, under Assumption 3.2, this mutual control relationship naturally extends to both D-optimality and E-optimality objectives. Specifically, their formulations simplify to
$\mathcal{L}_{D}(\mathbf{p}):=\det(\Sigma_{\mathbf{p}})=\prod_{i=1}^{d}\lambda_{i}(\mathbf{p}),\,\mathcal{L}_{E}(\mathbf{p}):=\lambda_{\max}(\Sigma_{\mathbf{p}}).$ (2)
It is evident that both $\mathcal{L}_{D}$ and $\mathcal{L}_{E}$ can be polynomially bounded above and below by $\mathcal{L}_{A}$ due to standard spectral inequalities. Consequently, through the control of $\mathcal{L}_{A}$, the flow-loss objective also polynomially controls $\mathcal{L}_{D}$ and $\mathcal{L}_{E}$ in both directions. Therefore, the flow-loss objective serves as a spectrum-aware surrogate that can also guide D-optimal and E-optimal sensor placement with bounded distortion.
@@ Appendix C Implementation Details
In this section, we provide the implementation details of our experiments, including benchmarks, metrics, and implementations. All the experiments are conducted based on PyTorch 2.1.0 Paszke2019PyTorchAI and on a A100 GPU server with 144 CPU cores.
@@ C.1 Benchmarks
We use a watertight 3D car mesh taken from the ShapeNet V1 chang2015shapenet dataset, which contains around 50k triangular facets. The aerodynamic pressure of the car surface is simulated using OpenFOAM jasak2009openfoam . As shown in Fig. 8(a), we create a virtual wind tunnel to simulate the airflow around a car. The width, length and height of the wind tunnel are set to 16 meters, 30 meters and 8 meters respectively. The computational mesh around the car surface is refined to a smaller mesh size for more accurate aerodynamics modeling, which is illustrated in Fig. 8(b).
@@ C.2 Metrics
where $\mathbf{u}$ denotes the reference (ground‑truth) field and $\hat{\mathbf{u}}$ represents the model reconstruction.
@@ C.3 Implementations
SSPOR manohar2018data
We utilize the official implementation of SSPOR, the PySensor package deSilva2021 to evaluate the reconstruction and sensor selection performance of SSPOR. SSPOR employs POD berkooz1993POD to express high-dimensional data as a linear combination of several orthonormal eigenmodes obtained from the singular value decomposition (SVD). Specifically, we employ the pysensors.SSPOR class with an SVD basis to obtain the targeted number of sensors from the training data. The selected sensors are then used to reconstruct dense fields from sparse observations either with SSPOR or our method.
VoronoiCNN fukami2021global
VoronoiCNN integrates Voronoi tessellation with convolutional neural networks (CNNs). However, the official implementation uses a simplistic CNN that struggles with complex and dynamic fields. We replace it with a more powerful U-Net ronneberger2015unet architecture, using 64 base feature channels, with 4 downsampling and upsampling layers and skip connections to enhance information flow. Due to its inability to handle irregular car aerodynamic meshes, experiments are conducted only on turbulent flow and sea temperature datasets under a unified setup.
Senseiver santos2023development
@@
DiffusionPDE huang2024diffusionpde
We borrow the model architectures of the original implementation of DiffusionPDE for the turbulent flow and sea temperature benchmarks, and employ a Transolver wu2024transolver as the backbone for the unstructured car aerodynamics benchmark. For inference, we also follow the original implementation configurations, using a total of 2,000 steps for better performance. The guidance weights we search in each benchmark are shown in Table 9.
@@ D.1 Comparsion with AdaLN-Zero
Here, we compare our sparse sensor incorporation strategy on turbulent flow with a widely used conditioning approach in diffusion models, namely the AdaLN-Zero block in DiT peebles2023dit . As shown in Table 12, replacing our design with AdaLN-Zero leads to a substantial degradation in performance, with reconstruction errors approaching 1. This is because sensor placement inherently captures localized information, whereas AdaLN-Zero aggregates all sensor data into a single global variable, which is insufficient for preserving the spatial specificity required in reconstruction tasks.
@@ D.3 Statistical Analysis
Furthermore, since our task is a reconstruction problem, and our reconstruction model essentially learns a conditional distribution, we report in the table above the mean and standard deviation of reconstruction loss under multiple samplings on the car aerodynamics benchmark, across varying numbers of sensors. As shown in Table 14 and 15, when the number of sensors exceeds 15, the standard deviation becomes relatively small, indicating stable performance. Moreover, regardless of the sensor count, once the placement is optimized, the reconstruction results exhibit consistent stability across multiple runs.
D.4 Additional Metric for Gradients
Table 16: Comparison on sea temperature benchmark. Relative L2 error of gradient fields is reported.
Sea Temperature (relative L2 of gradient field) #100 #50 #25 #15
Senseiver + random sensor placement 0.5456 0.5470 0.5501 0.5515
PhySense + random sensor placement 0.4877 0.4884 0.4890 0.4892
Relative promotion 11.87% 11.98% 12.51% 12.74%
In this section, we additionally evaluate the relative L2 error of the gradient fields on the sea temperature to better reflect whether the model captures meaningful physical fidelity. As shown in the Table 16, PhySense achieves significantly lower gradient error compared to the second-best model, Senseiver, demonstrating its superior spatial physical fidelity.
D.5 Uncertainty quantification
Figure 9: Visualization of uncertainty quantification for the flow-based reconstruction model on the turbulent flow benchmark. The sample variance is notably smaller than the physical field and error.
Flow-based reconstruction models explicitly learn the full conditional data distribution, which inherently facilitates uncertainty quantification. To complete this, we reconstruct each input three times, each with a distinct random noise initialization. Subsequently, we compute the pixel-wise sample mean and sample variance across these reconstructions. As illustrated in Fig. 9, the magnitude of the variance is orders of magnitude smaller than that of the sample mean and the reconstruction error. This indicates high confidence in our model, as it consistently produces similar reconstructions irrespective of the initial noise vector.
Appendix E Limitations and Future Work
While our current formulation assumes point-wise sensor observations, this may not fully capture the behavior of certain real-world sensors that integrate information over small spatial neighborhoods. For example, radar and precipitation sensors typically respond to local regions rather than single points. Incorporating regional observation into both the reconstruction model and the sensor placement optimization could further improve the effectiveness of our approach. As a future direction, we aim to extend both our model and theoretical framework from point-wise observations to region-based sensing, which better reflects the behavior of many real-world sensors.
Appendix F Boarder Impacts
We introduce PhySense, a synergistic two-stage framework for accurate physics sensing with theorectical guarantees. Extensive experiments validate its superior reconstruction accuracy and informative sensor placements. Beyond performance gains, our work offers a new perspective on the sensor placement problem from a deep learning standpoint. Thus, PhySense may inspire some future research in relevant domains. Moreover, our optimized placement could potentially inform how sensors are placed in more applications, enhancing their reconstruction accuracy.
Since our work is purely focused on algorithmic design for accurate physics sensing, there are no potential negative social impacts or ethical risks.